OpenStudy (anonymous):

64z^2 - 25y^2 how to factor??? HELP!

OpenStudy (cwrw238):

this is difference of 2 squares a^2 - b^2 = ( a + b)( a - b)

OpenStudy (cwrw238):

so if u put a^2 = 64z^2 and b^2 = 25y^2 you can hopefully see how to do it

OpenStudy (cwrw238):

u need to find square root of each

OpenStudy (anonymous):

im in a hurry , can u give me the answer to it man? its 8 and 5 but idk how to write it as a answer

OpenStudy (cwrw238):

a = 8z and b = 5y

OpenStudy (cwrw238):

just plug these into the a and b formula